OpenStudy (anonymous):

Write the expression as either sine, cos, or tan of a single angle sin23cos27-cos23sin27

OpenStudy (anonymous):

@bbnl1990

zepdrix (zepdrix):

Hey Alicia :) Remember your Sine Angle Difference Formula?\[\Large\rm \sin\left(\color{royalblue}{\alpha}-\color{orangered}{\beta}\right)=\sin\color{royalblue}{\alpha} \cos\color{orangered}{\beta}- \cos\color{royalblue}{\alpha}\sin\color{orangered}{\beta} \]

zepdrix (zepdrix):

Do you see how we can apply that to our problem here?\[\Large\rm \quad?\quad =\sin\color{royalblue}{23} \cos\color{orangered}{27}- \cos\color{royalblue}{23}\sin\color{orangered}{27}\]